Grade 12th passDifferential Calculus

f(x)+f(1-x)=2 , g(x)=f(x)-1 then g(x) is symmetrical about

praveen verma
9 Years agoGrade 12th pass

2 Answers

Abhi singhal
8 Years ago
f(x) + f(1-x) =2 and g(x)=f(X) – 1 then g(x) is symmetrical about 2 as we can see at 2 in graph there is a sharp edge and we know that at sharp edge the function is symmetrical
Vikas TU
8 Years ago
Dear Student,
Given Function as:
g(x)= f(x)-1
         So, f(x)= g(x)+1
      So, g(x)+1 + g(1-x)+1= 2
   Or,  g(x) = - g(1-x)
  Now , putting ½ - x  in place of x
   ,  g( ½-x) = - g( 1-1/2 + x)
                  = g( ½ + x)
   Therefore this is  symmetrical  about the point (1/2,0) 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)